# Rydberg superatoms: An artificial quantum system for quantum information processing and quantum optics

### v1

@@ II.2 Interaction potential between Rydberg atoms
where ${\rm H.c.}$ stands for the Hermitian conjugate and the real coefficient $C_{3}=\langle\uparrow\downarrow\left|\hat{d}_{1}^{+}\hat{d}_{2}^{-}|\downarrow\uparrow\right\rangle/(4\pi\epsilon_{0})$. This Hamiltonian was initially employed to examine the coherent dynamics of a spin excitation in a chain comprising three Rydberg atoms with $\theta=0$. Barredo et al. (2015); de Léséleuc et al. (2017) Subsequently, it was extended to simulate a topological model protected by symmetry in a zigzag configuration of Rydberg atoms, with the magic angle $\theta\approx 54.7^{\circ}$ governing the hopping direction along the next nearest neighbor atoms. de Léséleuc et al. (2019) Additionally, it was utilized to synthesize a density-dependent Peierls phase employing three Rydberg atoms arranged in an equilateral configuration with $\theta=\pi/2$. Lienhard et al. (2020) Furthermore, this model finds application in the simulation of energy transportSchönleber et al. (2015) and various generic spin chain models.Geier et al. (2021); Scholl et al. (2022); Chen et al. (2023a); Nishad et al. (2023); Chen et al. (2023b); Bornet et al. (2024); Li, Li, and Shao (2024)
The Förster resonance explores the resonant interaction between a pair of Rydberg states and other degenerate or quasi-degenerate energy levels under the dipole-dipole interaction described by Eq. (11). Even if a small energy defect exists, it can be mitigated by the differential Stark effect induced by external electric fields. Walker and Saffman (2005); Ryabtsev et al. (2010) The mathematical expression closely resembles Eq. (12), differing only in the composition of the bases. Experimental observations indicate that in the presence of a modest electric field ($F_{\rm res}\approx 33$ mV/cm), resonance occurs between the state $|dd\rangle$ and the symmetric state $|pf_{1}\rangle_{\rm s}=(|pf_{1}\rangle+|f_{1}p\rangle)/\sqrt{2}$ of two ${}^{87}$ Rb atoms, Ravets et al. (2014, 2015) where $|d\rangle=|59D_{3/2},m_{j}=3/2\rangle$, $|p\rangle=|61P_{1/2},m_{j}=1/2\rangle$ and $|f_{1}\rangle=|57F_{5/2},m_{j}=5/2\rangle$. Therefore, this interaction channel is effectively isolated from other interaction channels at an interatomic distance $R=9.1~{}\mu\rm m$ in the presence of a 3.3 G magnetic field. The corresponding Hamiltonian is given by
@@
In the absence of an external electric field, the energy difference defined by $\Delta_{r^{\prime},r^{\prime\prime}}=E_{r^{\prime}}+E_{r^{\prime\prime}}-2E_{r}$ of the Rydberg atom pair in the Förster resonance process greatly exceeds the strength of the dipole-dipole interaction when the atoms are at a significant distance from each other (perturbation theory is applicable). Here, $E_{r}$ represents the energy of the Rydberg state $|r\rangle$ of interest, while $E_{r^{\prime}}$ and $E_{r^{\prime\prime}}$ denote the energies of a pair of Rydberg states $|r^{\prime}\rangle$ and $|r^{\prime\prime}\rangle$, whose symmetric state is dipole coupled to the Rydberg pair state $|rr\rangle$. This leads to a multichannel-induced Stark shift in the energy levels of the atom pair, resulting in the interaction between the atoms manifesting as a classical van der Waals interaction,Walker and Saffman (2008); Saffman, Walker, and Mølmer (2010) represented by
@@
Here, $C_{6}$ is defined as $C_{6}=\sum_{r^{\prime}r^{\prime\prime}}|\langle r^{\prime}r^{\prime\prime}|\hat{V}_{int}|rr\rangle|^{2}/{\Delta_{r^{\prime},r^{\prime\prime}}}$, where in principle $|r^{\prime}r^{\prime\prime}\rangle$ encompasses all pair states dipole coupled to $|rr\rangle$. However, for practical computational efficiency, focusing on quantum states within a specific range of energy differences allows one to obtain sufficiently accurate values $C_{6}$. When the interatomic distance is not significantly large and perturbation theory is not applicable, it becomes crucial to perform the diagonalization of the Hamiltonian governing the pair-state interaction. Experimentally, the coefficients $|C_{6}|$ for $|nD_{3/2},m_{j}=3/2\rangle$ of the ${}^{87}$b atom with $n=53$, 62, and 82 are measured directly between two isolated single Rydberg atoms separated by a controlled distance of a few micrometers,Béguin et al. (2013) and the angular dependence of the effective interaction energy of the two $82D_{3/2}$ ($82S_{1/2}$) atoms is also demonstrated.Barredo et al. (2014) Recently, with the advancement of experimental techniques enabling the precise arrangement of a large number of atoms in defect-free programmable geometric configurations across various dimensions, the van der Waals-type Rydberg interaction has found applications in simulating large-scale Ising-type quantum spin models,Bernien et al. (2017); Lienhard et al. (2018); de Léséleuc et al. (2018a); Keesling et al. (2019); Ebadi et al. (2021); Scholl et al. (2021); Bharti et al. (2023); Kim et al. (2024) generating and manipulating Schrödinger cat states,Omran et al. (2019) designing frustrated magnets,Glaetzle et al. (2014, 2015) and exploring nonequilibrium physics.Bluvstein et al. (2021); Su et al. (2023); Shao et al. (2023); Zhao et al. (2024)
II.3 Emergence of Rydberg superatom
@@
In the fields of chemistry and atomic physics, the concept of a superatom is used to characterize a cluster of atoms that exhibits collective behavior and possesses characteristics distinct from its constituents.Knight et al. (1984); Inoshita, Ohnishi, and Oshiyama (1986); Khanna and Jena (1995) These superatoms possess unique electronic structures and reactivity,Luo and Castleman (2014); Reber and Khanna (2017) making them highly compelling for a wide range of applications in materials science and nanotechnology.Castleman Jr and Khanna (2009); Doud et al. (2020); Yu et al. (2023) Similarly, a Rydberg superatom denotes a collectively excited state of an ensemble of $N$ atoms, where the atoms exhibit such high correlation that only a single excitation is permitted, a phenomenon known as the Rydberg blockade.Jaksch et al. (2000) In a two-body system, this blockade effect can be achieved using either the Förster resonance or van der Waals interaction, as confirmed by experimental observations. Vogt et al. (2006); Urban et al. (2009); Gaëtan et al. (2009); Zeng et al. (2017) However, in a many-body system, the construction of a Rydberg superatom is typically facilitated by the Rydberg interaction in the form of a pairwise additive van der Waals potential.Heidemann et al. (2007, 2008); Dudin et al. (2012); Zeiher et al. (2015); Stiesdal et al. (2020) This is particularly true for the Rydberg-$S$ state, which possesses only one repulsive branch in its molecular potential, preventing a breakdown of the Rydberg blockade.Walker and Saffman (2005); Pohl and Berman (2009); Vermersch, Glaetzle, and Zoller (2015); Derevianko et al. (2015); de Léséleuc et al. (2018a) Here, we will provide a concise theoretical analysis of the mechanism and associated properties of Rydberg superatoms. Detailed explanations of the experimental procedures and their applications in quantum information and quantum optics will be discussed in the following sections.
@@
where a negative value of $C_{6}$ denotes repulsive van der Waals interactions, while a positive value signifies attractive interactions. At the same time, for the sake of simplification, we temporarily disregard the impact of atomic position and motion on the relative phase between individuals. Under the influence of Eq. (15), if all atoms in the ensemble are initially set in the ground state $|G\rangle=|g_{1},g_{2},...,g_{N}\rangle$, they will collectively transition to a symmetric single atom excitation state, assuming $C_{6}/R_{ij}^{6}\gg\hbar\Omega_{0}$. The critical interatomic distance denoted the Rydberg blockade radius and expressed as $r_{b}=[C_{6}/(\hbar\Omega_{0})]^{1/6}$,Heidemann et al. (2007); Browaeys and Lahaye (2020) plays a crucial role. Within this range, a single excitation of a Rydberg atom can impede the excitation of a second atom, defining the upper limit of the blockade range. When considering the collective excitation of two atoms, as shown in Fig. 2(a), the associated collective blockade radius is defined as $R_{b}=[C_{6}/(\hbar\sqrt{2}\Omega_{0})]^{1/6}$.Honer et al. (2011) Within this modified range, Rydberg atoms manifest a robust repulsive interaction, constraining excitation to a single excited symmetric state $|\psi_{+}\rangle=(|gr\rangle+|rg\rangle)/\sqrt{2}$ for the pair of atoms, assuming both atoms are driven with same Rabi frequency. An offset in Rabi frequencies of two atoms may freeze the dynamics in a single atom, inhibiting the creation of the collective single excited state.Srivastava, Niranjan, and Nath (2019) Expanding the system to the collective excitation of $N$ atoms, the collective blockade radius should be calculated as $R_{b}=[C_{6}/(\hbar\sqrt{N}\Omega_{0})]^{1/6}$. The corresponding collective state is represented by
$|R\rangle=\frac{1}{\sqrt{N}}\sum_{i=1}^{N}|g_{1},g_{2},g_{3},...,r_{i},...,g_{N}\rangle,$ (16)
@@
where all atoms share a single Rydberg excitation. Fig. 2(b) presents a schematic representation of a Rydberg superatom. This artificial atom comprises two effective levels: the ground state $|G\rangle$ and the collective excited state $|R\rangle$, connected by a coupling strength of $\sqrt{N}\Omega_{0}$. We note that the excited state of the Rydberg superatom is itself a multiparticle $W$ state with maximum entanglement.Dür, Vidal, and Cirac (2000) This entangled state holds significant relevance in quantum information processing.Cabello (2002); Horodecki et al. (2009); Kim et al. (2022) Recent experiments has demonstrated its robustness to both particle loss and external electric field noise, particularly as a logical qubit.Xu et al. (2021); Spong et al. (2021) Therefore, the emergent properties of a Rydberg superatom are derived from the collective behavior of its constituent atoms rather than from the properties of individual atoms alone. These properties encompass enhanced nonlinear optical responses and strong light-matter coupling, both of which hold promise for applications in quantum information processing and quantum optics. As the Rydberg superatom system progresses to the forefront of quantum science, it demonstrates the potential to revolutionize various technological domains, including quantum computing and advanced materials engineering. Ongoing research in this field explores various experimental techniques and theoretical models to comprehend and manipulate the behavior of Rydberg superatoms.
@@ III EXPERIMENTAL TECHNIQUES AND ADVANCES
III.1 Coherent laser excitation of Rydberg atom
@@
An essential step in Rydberg experiments is the preparation of Rydberg atoms. The utilization of lasers offers high-resolution, coherent, and efficient excitation of Rydberg states. After the invention of tunable dye lasers, Rydberg laser spectroscopy with resolution $\sim 0.1~{}\rm nm$ was demonstrated in 1973.Bradley et al. (1973); Stebbings and Dunning (1983) The spectral resolution was improved to $\sim$ 10 GHz in 1979. In 2003, Rydberg Autler-Townes spectra with MHz resolution were studied with narrow linewidth lasers. Teo et al. (2003)
@@
Figure 3: Different Rydberg laser excitation schemes for ${}^{87}$b atom. 297 nm UV laser can couple the ground state $5S$ and the Rydberg state $nP$. 780 nm (420 nm) and 480 nm (1012 nm) lasers can excite atoms from $5S$ to $nS$ or $nD$ via an intermediate state $5P$ $(6P)$ with intermediate detuning $\Delta_{e}$. Due to the finite lifetime of the intermediate states, atoms in these states decay with a rate of $\Gamma_{e}$, which is one of the main sources of error in the two-photon excitation scheme. Using the intermediate states $5P$ and $6S$, atoms can be excited to the $nP$ Rydberg states using the 780 nm, 1367 nm, and 743 nm lasers.
@@
Figure 4: Laser setup of 480 nm and 780 nm lasers used in ${}^{87}$b Rydberg excitation. The 780   nm and 960   nm lasers are obtained from external cavity diode lasers (ECDLs). Both lasers are frequency locked to a high finesse, ultra-stable cavity using the Pound-Drever-Hall (PDH) locking technique, which generates “servo bump” due to the limited stabilization bandwidth. Frequency doubling of the power-amplified (using the tapered amplifier) 960   nm laser in the second harmonic generator (SHG) generates blue light at 480 nm. The 780 nm and 480 nm laser beams excite the atoms in a counterpropagating configuration to minimize the Doppler effect.
@@
The laser fields used for Rydberg excitation must be intense enough to guarantee a fast excitation within the Rydberg lifetime and coherence time. The natural lifetimes of high-lying Rydberg states are mainly limited by spontaneous emission and black-body radiation (BBR), and can last more than hundreds of microseconds. However, the Rydberg-ground coherence times are usually about a few to tens of microseconds, limited by many decoherence channels, such as the Doppler effect, density-dependent phase shift, fluctuating background electric fields, etc. Therefore, the Rabi frequencies of the excitation laser have to be large, usually on the orders of MHz, to ensure coherent excitations. Such large Rabi frequencies require intense light sources for highly excited Rydberg states since the dipole matrix elements between the ground and Rydberg states decrease with the principal quantum number as $n^{*-1.5}$. In Rydberg tweezer array experiments, it has been shown that a larger Rabi frequency is crucial for high-fidelity operations within the finite ground-Rydberg coherence.Madjarov et al. (2020) In a Rydberg superatom containing $N$ atoms, the effective ground-Rydberg Rabi frequency is improved by a factor of $\sqrt{N}$ due to collective enhancement effects.
@@
Given the considerable energy gaps between the ground and Rydberg states, direct ground-to-Rydberg transition frequencies typically fall within the ultra-violet (UV) range.Wang et al. (2017) Taking Rubidium atoms as an example, lasers with $\sim$ 297 nm wavelength can promote atoms from the $5S$ ground state to the $nP$ Rydberg states.Thoumany et al. (2009) Coherent manipulation between the ground and Rydberg states using a UV laser was demonstrated in Fig. 3, see Ref. Hankin et al., 2014. The single photon UV scheme is ideal for realizing the Rydberg dressing and has also been used in the Rydberg-ground state quantum memory.Li and Kuzmich (2016)
@@
To avoid the use of UV Rydberg lasers, a commonly adopted scheme involves an intermediate state and use two-photon Rydberg excitation, as shown in Fig. 3. With the 780 nm - 480 nm (420 nm - 1012 nm) two-photon scheme, the ground state Rb atom can be excited to the $nS$ or $nD$ Rydberg states. Fig. 4 shows the laser setup for the 780 nm - 480 nm scheme. The 780 nm light field can be easily obtained from a commercially available external cavity diode laser (ECDL), while the 480 nm light comes from a second-harmonic-generator (SHG) seeded by power-amplified 960 nm laser light. To achieve a strong blockade effect, the laser linewidth must be narrow enough compared to the Rydberg interactions. Therefore, both 780 nm and 960 nm ECDLs are frequency stabilized to an ultralow expansion cavity with high finesse to guarantee narrow linewidth. However, the Pound–Drever–Hall (PDH) laser locking technique leads to large “servo bumps” outside the stabilization bandwidth, which could cause severe decoherence during laser excitation. To suppress this noise, cavity-filtering Levine et al. (2018) or active feedforward techniquesLi et al. (2022a) can be used.
@@
The Rydberg excitation fidelity is affected by various decoherence and loss mechanisms. For the single-photon scheme, a major source of errors in coherent excitation between ground and Rydberg states is the Doppler effect due to atomic motions, which results in a random detuning $\delta=kv$, where $v$ is the velocity of the atom along the wave vector of the laser beam $k$. Limited by the available UV laser power and the small dipole moment between the ground and Rydberg states, the Rabi frequency of the single-photon excitation $\Omega_{297}$ is on the order of a few MHz. As shown in Fig. 5(a), the ground-Rydberg states $\pi$-pulse excitation error exceeds a few percent for an atomic temperature of ${50}~{}{\mu}$K and $\Omega_{297}\sim$ MHz. Infidelity caused by the Doppler effect can be suppressed by lowering the temperature or increasing the Rabi frequency.
Compared to the UV laser in single-photon excitation, it is easier to achieve high power for the visible and near-infrared lasers used in two-photon schemes. Therefore, the Rabi frequency of the ground-Rydberg state $\Omega=\Omega_{1}\Omega_{2}/2\Delta_{e}$ in two-photon excitation can reach tens or even hundreds of MHz, where $\Omega_{1}$ $(\Omega_{2})$ is the Rabi frequency for the ground-intermediate (intermediate-Rydberg) transition, and $\Delta_{e}$ is the intermediate-state detuning, as shown in Fig. 3.
In the two-photon scheme, if a counterpropagating excitation protocol is employed, the Doppler effect will be less severe than in the single-photon case. As shown in Fig. 4, by counterpropagating the two excitation beams, the Doppler effect can be minimized. It is also obvious that the 780 nm - 480 nm excitation scheme has a smaller Doppler-induced error than that of the 420 nm - 1012 nm scheme. However, regardless of the excitation approach used, achieving a Doppler-induced error well below 1% is achievable by either lowering the atomic temperature or applying a strong laser (large Rabi frequency).

### v2

@@ II.2 Interaction potential between Rydberg atoms
where ${\rm H.c.}$ stands for the Hermitian conjugate and the real coefficient $C_{3}=\langle\uparrow\downarrow\left|\hat{d}_{1}^{+}\hat{d}_{2}^{-}|\downarrow\uparrow\right\rangle/(4\pi\epsilon_{0})$. This Hamiltonian was initially employed to examine the coherent dynamics of a spin excitation in a chain comprising three Rydberg atoms with $\theta=0$. Barredo et al. (2015); de Léséleuc et al. (2017) Subsequently, it was extended to simulate a topological model protected by symmetry in a zigzag configuration of Rydberg atoms, with the magic angle $\theta=\arccos(1/\sqrt{3})$ governing the hopping direction along the next nearest neighbor atoms. de Léséleuc et al. (2019) Additionally, it was utilized to synthesize a density-dependent Peierls phase employing three Rydberg atoms arranged in an equilateral configuration with $\theta=\pi/2$. Lienhard et al. (2020) Furthermore, this model finds application in the simulation of energy transportSchönleber et al. (2015) and various generic spin chain models.Geier et al. (2021); Scholl et al. (2022); Chen et al. (2023a); Nishad et al. (2023); Chen et al. (2023b); Bornet et al. (2024); Li, Li, and Shao (2024)
When the electronic energy of a pair of Rydberg atoms is (nearly) degenerate with energies of other Rydberg pair states, the dipole-dipole interaction can be strong due to the so-called Föster resonance. Even if a small energy defect exists, it can be mitigated by the differential Stark effect induced by external electric fields. Walker and Saffman (2005); Ryabtsev et al. (2010) The mathematical expression closely resembles Eq. (12), differing only in the composition of the bases. Experimental observations indicate that in the presence of a modest electric field ($F_{\rm res}\approx 33$ mV/cm), resonance occurs between the state $|dd\rangle$ and the symmetric state $|pf_{1}\rangle_{\rm s}=(|pf_{1}\rangle+|f_{1}p\rangle)/\sqrt{2}$ of two 87 Rb atoms, Ravets et al. (2014, 2015) where $|d\rangle=|59D_{3/2},m_{j}=3/2\rangle$, $|p\rangle=|61P_{1/2},m_{j}=1/2\rangle$ and $|f_{1}\rangle=|57F_{5/2},m_{j}=5/2\rangle$. Therefore, this interaction channel is effectively isolated from other interaction channels at an interatomic distance $R=9.1~{}\mu\rm m$ in the presence of a 3.3 G magnetic field. The corresponding Hamiltonian is given by
@@
In the absence of an external electric field, the energy difference defined by $\Delta_{r^{\prime},r^{\prime\prime}}=E_{r^{\prime}}+E_{r^{\prime\prime}}-2E_{r}$ of the Rydberg atom pair in the Förster resonance process greatly exceeds the strength of the dipole-dipole interaction when the atoms are at a significant distance from each other (perturbation theory is applicable). Here, $E_{r}$ represents the energy of the Rydberg state $|r\rangle$ of interest, while $E_{r^{\prime}}$ and $E_{r^{\prime\prime}}$ denote the energies of a pair of Rydberg states $|r^{\prime}\rangle$ and $|r^{\prime\prime}\rangle$, whose symmetric state is dipole coupled to the Rydberg pair state $|rr\rangle$. This leads to a multichannel-induced Stark shift in the energy levels of the atom pair, resulting in the interaction between the atoms manifesting as a classical vdw interaction,Walker and Saffman (2008); Saffman, Walker, and Mølmer (2010) represented by
@@
Here, $C_{6}$ is defined as $C_{6}=\sum_{r^{\prime}r^{\prime\prime}}|\langle r^{\prime}r^{\prime\prime}|\hat{V}_{int}|rr\rangle|^{2}/{\Delta_{r^{\prime},r^{\prime\prime}}}$, where in principle $|r^{\prime}r^{\prime\prime}\rangle$ encompasses all pair states dipole coupled to $|rr\rangle$. However, for practical computational efficiency, focusing on quantum states within a specific range of energy differences allows one to obtain sufficiently accurate values of $C_{6}$. When the interatomic distance is not significantly large and perturbation theory is not applicable, it becomes crucial to perform the diagonalization of the Hamiltonian governing the pair-state interaction. Experimentally, the $|C_{6}|$ coefficients for $|nD_{3/2},m_{j}=3/2\rangle$ of the 87Rb atom with $n=53$, 62, and 82 are measured directly between two isolated single Rydberg atoms separated by a controlled distance of a few micrometers,Béguin et al. (2013) and the angular dependence of the effective interaction energy of the two $82D_{3/2}$ ($82S_{1/2}$) atoms is also demonstrated.Barredo et al. (2014) Recently, with the advancement of experimental techniques enabling the precise arrangement of a large number of atoms in defect-free programmable geometric configurations across various dimensions, the vdw-type Rydberg interaction has found applications in simulating large-scale Ising-type quantum spin models,Bernien et al. (2017); Lienhard et al. (2018); de Léséleuc et al. (2018a); Keesling et al. (2019); Ebadi et al. (2021); Scholl et al. (2021); Bharti et al. (2023); Kim et al. (2024) generating and manipulating Schrödinger cat states,Omran et al. (2019) designing frustrated magnets,Glaetzle et al. (2014, 2015) and exploring nonequilibrium physics.Ding et al. (2020); Bluvstein et al. (2021); Su et al. (2023); Shao et al. (2023); Zhao et al. (2024)
II.3 Emergence of Rydberg superatoms
@@
In the fields of chemistry and atomic physics, the concept of a superatom is used to characterize a cluster of atoms that exhibits collective behavior and possesses characteristics distinct from its constituents.Knight et al. (1984); Inoshita, Ohnishi, and Oshiyama (1986); Khanna and Jena (1995) These superatoms possess unique electronic structures and reactivity,Luo and Castleman (2014); Reber and Khanna (2017) making them highly compelling for a wide range of applications in materials science and nanotechnology.Castleman Jr and Khanna (2009); Doud et al. (2020); Yu et al. (2023) Similarly, a Rydberg superatom denotes a collectively excited state of an ensemble of $N$ atoms, where the Rydberg blockade inhibits the simultaneous excitation of multiple Rydberg atoms in the ensemble.Jaksch et al. (2000) In a two-body system, this blockade effect can be achieved using either the Förster resonance or vdw interaction, as confirmed by experimental observations. Vogt et al. (2006); Urban et al. (2009); Gaëtan et al. (2009); Zeng et al. (2017) However, in a many-body system, the construction of a Rydberg superatom is typically facilitated by the Rydberg interaction in the form of a pairwise additive vdw potential.Heidemann et al. (2007, 2008); Dudin et al. (2012); Zeiher et al. (2015); Stiesdal et al. (2020) This is particularly true for the Rydberg-$S$ state, which possesses only one repulsive branch in its molecular potential, preventing a breakdown of the Rydberg blockade.Walker and Saffman (2005); Pohl and Berman (2009); Vermersch, Glaetzle, and Zoller (2015); Derevianko et al. (2015); de Léséleuc et al. (2018a) Here, we will provide a concise theoretical analysis of the mechanism and associated properties of Rydberg superatoms. Detailed explanations of the experimental procedures and their applications in quantum information and quantum optics will be discussed in the following sections.
@@
where a negative value of $C_{6}$ denotes repulsive vdw interactions, while a positive value signifies attractive interactions. At the same time, for the sake of simplification, we temporarily disregard the impact of atomic position and motion on the relative phase between individuals. Under the influence of Eq. (15), if all atoms in the ensemble are initially set in the ground state $|G\rangle=|g_{1},g_{2},...,g_{N}\rangle$, they will collectively transition to a symmetric single atom excitation state, assuming $C_{6}/R_{ij}^{6}\gg\hbar\Omega_{0}$. When examining the simple case of the collective excitation of two atoms, as illustrated in Fig. 2(a), the associated collective blockade radius is defined as $R_{b}=[C_{6}/(\hbar\sqrt{2}\Omega_{0})]^{1/6}$,Honer et al. (2011) where the interaction energy equals the linewidth of excitation dominated by the power broadening $\sqrt{2}\Omega_{0}$. Within this range, Rydberg atoms manifest a robust repulsive interaction, constraining excitation to a single excited symmetric state $|\psi_{+}\rangle=(|gr\rangle+|rg\rangle)/\sqrt{2}$ for the pair of atoms, assuming both atoms are driven with the same Rabi frequency. An offset in Rabi frequencies of two atoms may freeze the dynamics in a single atom, inhibiting the creation of the collective single excited state.Srivastava, Niranjan, and Nath (2019) Expanding the system to the collective excitation of $N$ atoms, the collective blockade radius should be calculated as $R_{b}=[C_{6}/(\hbar\sqrt{N}\Omega_{0})]^{1/6}$, and the corresponding collective state is represented by
$|R\rangle=\frac{1}{\sqrt{N}}\sum_{j=1}^{N}|g_{1},g_{2},...,r_{j},...,g_{N}\rangle,$ (16)
@@
where all atoms share a single Rydberg excitation. This form of collective excitation also exists in the single-photon superradiance model, where identical ground state atoms interact with a single photon.Scully et al. (2006); Svidzinsky, Chang, and Scully (2008); Scully (2009, 2015); Wang et al. (2015) Figure 2(b) presents a schematic representation of a Rydberg superatom, which is collectively coupled to the ground state $|G\rangle$ with a coupling strength of $\sqrt{N}\Omega_{0}$. We note that the Rydberg superatom is itself a multiparticle $W$ state with maximum entanglement.Dür, Vidal, and Cirac (2000) This entangled state holds significant relevance in quantum information processing.Cabello (2002); Horodecki et al. (2009); Kim et al. (2022) Recent experiments has demonstrated its robustness to both particle loss and external electric field noise, particularly as a logical qubit.Xu et al. (2021); Spong et al. (2021) Therefore, the emergent properties of a Rydberg superatom are derived from the collective behavior of its constituent atoms rather than from the properties of individual atoms alone. These properties encompass enhanced nonlinear optical responses and strong light-matter coupling, both of which hold promise for applications in quantum information processing and quantum optics. As the Rydberg superatom system progresses to the forefront of quantum science, it demonstrates the potential to revolutionize various technological domains, including quantum computing and advanced materials engineering. One important theme in the ongoing research is to develop experimental techniques and explore theoretical models to understand and manipulate properties and dynamics of Rydberg superatoms.
@@ III EXPERIMENTAL TECHNIQUES AND ADVANCES
III.1 Coherent laser excitation of Rydberg atoms
@@
An essential step in Rydberg experiments is the preparation of Rydberg atoms. The utilization of lasers offers high-resolution, coherent, and efficient excitation of Rydberg states. After the invention of tunable dye lasers, Rydberg laser spectroscopy with resolution $\sim 0.1\rm~{}nm$ was demonstrated in 1973.Bradley et al. (1973); Stebbings and Dunning (1983) The spectral resolution was improved to $\sim$ 10 GHz in 1979.Zimmerman et al. (1979) In 2003, Rydberg Autler-Townes spectra with MHz resolution were studied with narrow linewidth lasers. Teo et al. (2003)
@@
Figure 3: Different Rydberg laser excitation schemes for 87Rb atom. 297 nm UV laser can couple the ground state $5S$ and the Rydberg state $nP$. 780 nm (420 nm) and 480 nm (1012 nm) lasers can excite atoms from $5S$ to $nS$ or $nD$ via an intermediate state $5P$ $(6P)$ with intermediate detuning $\Delta_{e}$. Due to the finite lifetime of the intermediate states, atoms in these states decay with a rate of $\Gamma_{e}$, which is one of the main sources of error in the two-photon excitation scheme. Using the intermediate states $5P$ and $6S$, atoms can be excited to the $nP$ Rydberg states using the 780 nm, 1367 nm, and 743 nm lasers.
@@
Figure 4: Laser setup of 480 nm and 780 nm lasers used in 87Rb Rydberg excitation. The 780 nm and 960 nm lasers are obtained from external cavity diode lasers (ECDLs). Both lasers are frequency locked to a high finesse, ultra-stable cavity using the Pound-Drever-Hall (PDH) locking technique, which generates “servo bump” due to the limited stabilization bandwidth. Frequency doubling of the power-amplified (using the tapered amplifier) 960 nm laser in the second harmonic generator (SHG) generates blue light at 480 nm. The 780 nm and 480 nm laser beams excite the atoms in a counterpropagating configuration to minimize the Doppler effect.
@@
The laser fields used for Rydberg excitation must be intense enough to guarantee a fast excitation within the Rydberg lifetime and coherence time. The natural lifetimes of high-lying Rydberg states are mainly limited by spontaneous emission and black-body radiation, and can last more than hundreds of microseconds. However, the ground-Rydberg coherence times are usually about a few to tens of microseconds, limited by many decoherence channels, such as the Doppler effect, density-dependent phase shift, fluctuating background electric fields, etc. Therefore, the Rabi frequencies of the excitation laser have to be large, usually on the orders of MHz, to ensure coherent excitations. Such large Rabi frequencies require intense light sources for highly excited Rydberg states since the dipole matrix elements between the ground and Rydberg states decrease with the principal quantum number as $n^{*-1.5}$. In Rydberg tweezer array experiments, it has been shown that a larger Rabi frequency is crucial for high-fidelity operations within the finite ground-Rydberg coherence.Madjarov et al. (2020) In a Rydberg superatom containing $N$ atoms, the effective ground-Rydberg Rabi frequency is improved by a factor of $\sqrt{N}$ due to collective enhancement effects.
@@
Given the considerable energy gaps between the ground and Rydberg states, direct ground-to-Rydberg transition frequencies typically fall within the ultra-violet (UV) range.Wang et al. (2017) Taking Rubidium atoms as an example, lasers with $\sim$ 297 nm wavelength can excite atoms from the $5S$ ground state to the $nP$ Rydberg states.Thoumany et al. (2009) Coherent manipulation between the ground and Rydberg states using a UV laser was demonstrated in Fig. 3, see Ref. Hankin et al., 2014. The single photon UV scheme is ideal for realizing the Rydberg dressing and has also been used in the ground-Rydberg state quantum memory.Li and Kuzmich (2016)
@@
The choice of the excitation scheme depends on the experimental aim. A commonly adopted scheme involves an intermediate state and uses two-photon Rydberg excitation, which can benefit from electromagnetically induced transparency (EIT). With the 780 nm-480 nm (420 nm-1012 nm) two-photon scheme, the ground state Rb atom can be excited to the $nS$ or $nD$ Rydberg states, as shown in Fig. 3. Figure 4 shows the laser setup for the 780 nm-480 nm scheme. The 780 nm light field can be easily obtained from a commercially available external cavity diode laser (ECDL), while the 480 nm light comes from a second-harmonic-generator (SHG) seeded by power-amplified 960 nm laser light. To achieve a strong blockade effect, the laser linewidth must be narrow enough compared to the Rydberg interactions. Therefore, both 780 nm and 960 nm ECDLs are frequency stabilized to an ultralow expansion cavity with high finesse to guarantee narrow linewidth. However, the Pound-Drever-Hall (PDH) laser locking technique leads to large “servo bumps” outside the stabilization bandwidth, which could cause severe decoherence during laser excitation.
@@
The Rydberg excitation fidelity is affected by various decoherence and loss mechanisms. For the single-photon scheme, a major source of errors in coherent excitation between ground and Rydberg states is the Doppler effect due to atomic motions, which results in a random detuning $\delta=kv$, where $v$ is the velocity of the atom along the wave vector of the laser beam $k$. Limited by the available UV laser power and the small dipole moment between the ground and Rydberg states, the Rabi frequency of the single-photon excitation $\Omega_{\rm UV}$ is on the order of a few MHz. As shown in Fig. 5(a), the ground-Rydberg states $\pi$-pulse excitation error exceeds a few percent for an atomic temperature of ${50}~{}{\mu}$K and $\Omega_{\rm UV}\sim$ MHz. Infidelity caused by the Doppler effect can be suppressed by lowering the temperature or increasing the Rabi frequency.
Compared to the UV laser in single-photon excitation, it is easier to achieve high power for the visible and near-infrared lasers used in two-photon schemes. Therefore, the Rabi frequency of the ground-Rydberg state $\Omega=\Omega_{1}\Omega_{2}/(2\Delta_{e})$ in two-photon excitation can reach tens or even hundreds of MHz, where $\Omega_{1}$ $(\Omega_{2})$ is the Rabi frequency for the ground-intermediate (intermediate-Rydberg) transition, and $\Delta_{e}$ is the intermediate-state detuning, as shown in Fig. 3.
In the two-photon scheme, if a counterpropagating excitation protocol is employed, the Doppler effect will be less severe than in the single-photon case. As shown in Fig. 4, by using counterpropagating excitation beams, the Doppler effect can be minimized.Levine et al. (2018) It is also obvious that the 780 nm-480 nm excitation scheme has a smaller Doppler-induced error than that of the 420 nm-1012 nm scheme. However, regardless of the excitation approach used, achieving a Doppler-induced error well below 1% is achievable by either lowering the atomic temperature or applying a strong laser (large Rabi frequency).